4y/5 - 1/4 + y/10 = 7/2
Answer by ewatrrr(24785)

 
Hi
4y/5 - 1/4 + y/10 = 7/2 |Multiplying thru by 20 so as all denominators = 1
16y - 5 + 2y = 70
18y = 75
y = 75/18 = 25/6 or 4 1/6